# Thermal and Quantum Phase Transitions of the $\phi^{4}$ Model

### v1

@@ IV Thermal RG study of the $\phi^{4}$ model in lower dimensions
Figure 2: Thermal RG flow diagram of the $\phi^{4}$ model in $d=3$ dimensions based on flow equations (IV) and (IV). Circles show how the position of the Wilson-Fisher fixed point changes with $\tau$.
@@
Figure 3: Thermal RG flow diagram of the $\phi^{4}$ model in $d=3$ dimensions with $\tau=1$ based on the RG flow equations (IV) and (IV). Black cross denotes an initial condition which lies on the separatrix. For $\tau<1$ the RG trajectory from that starting point runs into the broken (low-temperature) phase and for $\tau>1$ the RG trajectory ends up in the symmetric (high-temperature) phase.
@@ V QPT-CPT diagram of the $\phi^{4}$ model
These conclusions can be drawn without performing any explicit calculations on the QPT-CPT diagram of the $\phi^{4}$ model. In other words, one can assume the above statements which can be used to test the thermal RG equations (6) and (II.2) suggested by us.
The QPT-CPT diagram of the three-dimensional $\phi^{4}$ model is constructed by the numerical solution of the thermal RG equations (6) and (II.2), see the solid (Litim cutoff) and the dashed (shapr-cutoff) lines of Fig. 5.
Figure 5: QPT-CPT diagram of the $\phi^{4}$ model in $d=3$ dimensions based on the thermal RG equations (6) and (II.2). The solid (dashed) line stands for results obtained by the Litim (sharp-cutoff) regulator. The inset shows the critical exponent $\nu$ as a function of $\tau$ which is related to the temperature. The shaded areas stand for $\tau<1$.
@@ VI Conclusions
It is important to note that previous studies based on the finite-temperature FRG method had to discuss findings with special emphasis on the interplay of the temperature $T$, and the running RG scale $k$. Above Eq. (58) of thermal_rg_phi4_volume one finds for example, the RG flow equation of the infinite volume Stefan–Boltzmann pressure $\partial_{k}p(T)=-k^{4}/12\pi^{2}[\coth(k/2T)-1]$. To draw conclusions on, one has to discuss how the running RG scale $k$ is related to the temperature which is fixed to the UV cutoff, i.e., $T=\tau\Lambda$. For large cutoff scales, $k>>T$ the flow decays exponentially with $\exp(-k/T)$ which is given by Eq. (58) of thermal_rg_phi4_volume . By using our suggestions where the temperature parameter is related to the running scale $T=\tau k$, and the dimensionless parameter $\tau$ plays the role of the temperature, this flow equation is simplified as $\partial_{k}p(\tau)=-k^{4}/12\pi^{2}[\coth(1/2\tau)-1]$ and conclusions can be drawn based on the change of the dimensionless parameter $\tau$ only. Thus, the modified thermal RG method of this work finds application in a natural way in any previous thermal RG studies and simplifies the discussion.
@@ Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” are gratefully acknowledged.

### v2

@@ V Thermal RG study of the $\phi^{4}$ model in lower dimensions
Figure 2: Thermal RG flow diagram of the $\phi^{4}$ model in $d=3$ dimensions based on flow equations (V) and (V). Circles show how the position of the WF fixed point changes with $\tau$.
@@
Figure 3: Thermal RG flow diagram of the $\phi^{4}$ model in $d=3$ dimensions with $\tau=1$ based on the RG flow equations (V) and (V). Black cross denotes an initial condition which lies on the separatrix. For $\tau<1$ the RG trajectory from that starting point runs into the broken (low-temperature) phase and for $\tau>1$ the RG trajectory ends up in the symmetric (high-temperature) phase.
@@ VI QPT-CPT diagram of the $\phi^{4}$ model
These conclusions can be drawn without performing any explicit calculations on the QPT-CPT diagram of the $\phi^{4}$ model. In other words, one can assume the above statements which can be used to test the thermal RG equations (6) and (III.2) suggested by us.
The QPT-CPT diagram of the three-dimensional $\phi^{4}$ model is constructed by the numerical solution of the thermal RG equations (6) and (III.2), see the solid (Litim cutoff) and the dashed (shapr-cutoff) lines of Fig. 5.
Figure 5: QPT-CPT diagram of the $\phi^{4}$ model in $d=3$ dimensions based on the thermal RG equations (6) and (III.2). The solid (dashed) line stands for results obtained by the Litim (sharp-cutoff) regulator. The inset shows the critical exponent $\nu$ as a function of $\tau$ which is related to the temperature. The shaded areas stand for $\tau<1$.
@@ VII Conclusions
It is important to note that previous studies based on the finite-temperature FRG method had to discuss findings with special emphasis on the interplay of the temperature $T$, and the running RG scale $k$. Above Eq. (58) of thermal_rg_phi4_volume one finds for example, the RG flow equation of the infinite volume Stefan–Boltzmann pressure $\partial_{k}P(T)=-k^{4}/12\pi^{2}[\coth(k/2T)-1]$. To draw conclusions on, one has to discuss how the running RG scale $k$ is related to the temperature which is fixed to the UV cutoff, i.e., $T=\tau\Lambda$. For large cutoff scales, $k>>T$ the flow decays exponentially with $\exp(-k/T)$ which is given by Eq. (58) of thermal_rg_phi4_volume . By using our suggestions where the temperature parameter is related to the running scale $T=\tau k$, and the dimensionless parameter $\tau$ plays the role of the temperature, this flow equation is simplified as $\partial_{k}P(\tau)=-k^{4}/12\pi^{2}[\coth(1/2\tau)-1]$ and conclusions can be drawn based on the change of the dimensionless parameter $\tau$ only. Thus, the modified thermal RG method of this work finds application in a natural way in all previous thermal RG studies of quantum phase transitions and simplifies the discussion.
@@ Acknowledgements
The support for the CNR/MTA Italy-Hungary 2023-2025 Joint Project ”Effects of strong correlations in interacting many-body systems and quantum circuits” is gratefully acknowledged.